# On the use of the Gram matrix for multivariate functional principal components analysis

### v1

@@ 3.4 On centering and reducing
For conducting an MFPCA, the features are usually assumed centred (Happ and Greven,, 2018). Prothero et al., (2023) give a complete overview of centering in the context of FDA. Here, we comment on the geometric interpretation of centering in this context and compare with the multivariate scalar case. We focus on the usual centering in FDA, namely $X_{n}^{(p)}(t_{p})-\mu^{(p)}(t_{p}),\leavevmode\nobreak\ t_{p}\in\mathcal{T}_{p}$ (refered as object centering in Prothero et al., (2023)). The geometric interpretation of object centering is the same if we refer to the observation space $\mathcal{H}$ or the feature space $\mathbb{R}^{N}$. Within the space $\mathcal{H}$ (resp. $\mathbb{R}^{N}$), centering is interpreted as translating the centre of gravity of the clouds, $\mathrm{G}_{\!\mu}$ (resp. $\mathsf{G}_{\!\mu}$), to the origin point $\mathrm{O}_{\!\mathcal{H}}$ (resp. $\mathsf{O}_{\!\mathbb{R}}$) of the space $\mathcal{H}$ (resp. $\mathbb{R}^{N}$). This transformation, being a translation, does not change the shape of the cloud $\mathcal{C}_{\!P}$ (resp. $\mathcal{C}_{\!N}$). The interpretation is the same as for the centering in the multivariate scalar data context within their observation space.
Concerning the standardization of the data, there are two main proposals in the literature. Happ and Greven, (2018) propose to weight each component $p$ by
@@
This standardization is coherent with the derivation of the total inertia of the observation space using the usual distance in $\mathcal{H}$. Chiou et al., (2014) propose to standardize each component $p$ of the data using the function
$w_{p}(t_{p})=\left(\operatorname{Var}X^{(p)}(t_{p})\right)^{-1/2},\quad t_{p}\in\mathcal{T}_{p}.$ (27)
This corresponds to a standardization of the curves by the standard deviation of the component at each sampling point. The standard deviation curve is estimated as the square root of the diagonal of the covariance function estimates, obtained using a local linear smoother of the pooled data. For each functional feature $p$, this standardization mimics the standardization used for principal components analysis if the number of (scalar) features is infinite. Considering the duality diagram and the total inertia of the clouds with respect to the distance $d_{\Gamma}$, we propose to weight each component $p$ by
$w_{p}=\left(\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}\right)^{-1/2}.$ (28)
The total inertia of $\mathcal{C}_{\!P}$ (and $\mathcal{C}_{\!N}$) will thus be equal to the number of components.
@@ 4 Multivariate functional principal components analysis
where $\mathfrak{c}_{k}=\langle\!\langle X-\mu,\phi_{k}\rangle\!\rangle$ are the projections of the centered curve onto the eigenfunctions. We have that $\mathbb{E}(\mathfrak{c}_{k})=0$, $\mathbb{E}(\mathfrak{c}_{k}^{2})=\lambda_{k}$ and $\mathbb{E}(\mathfrak{c}_{k}\mathfrak{c}_{k^{\prime}})=0$ for $k\neq k^{\prime}$. Note that the coefficients $\mathfrak{c}_{k}$ are scalar random variables while the multivariate functions $\phi_{k}$ are vectors of functions. Let us call $\Phi$ the multivariate functional principal component analysis basis. In practice, we use a truncated version of the Karhunen-Loève expansion (30) as the eigenvalues $\lambda_{k}$, and hence the contribution of $\mathfrak{c}_{k}$ to (30), becomes negligible as $k$ goes to infinity. Let
@@
be the truncated Karhunen-Loève expansion of the $p$th feature of the process $X$. For each $p$, the set $\{\varphi_{k}\}_{1\leq k\leq K_{p}}$ is a basis of univariate functions in $\mathcal{L}^{2}\left(\mathcal{T}_{p}\right)$, whose elements are not the components of the multivariate functions $\phi_{k}$.
@@ 4.1 Diagonalization of the covariance operator
[…]
@@
$\mathfrak{c}_{nk}=\mathcal{Z}_{{n,\cdot}}\boldsymbol{v}_{k},\quad 1\leq n\leq N,\quad 1\leq k\leq K_{+}.$
We refer the reader to Happ and Greven, (2018) for the derivation of the eigencomponents of the covariance operator if the curves are expanded in a general basis of functions.
@@ 4.2 Diagonalization of the inner product matrix
We can use the duality relation between row and column spaces of a data matrix to estimate the eigencomponents of the covariance operator. Consider the inner-product matrix $\mathbf{M}$, with entries defined in (6) and assuming that all observations are equally weighted, i.e., for all $n=1,\dots,N$, $\pi_{n}=1/N$. Let $\{l_{k}\}_{1\leq k\leq N}$ such that $l_{1}\geq\dots\geq l_{N}\geq 0$ be the set of eigenvalues and $\{v_{k}\}_{1\leq k\leq N}$ be the set of eigenvectors of the matrix $\mathbf{M}$. The relationship between all nonzero eigenvalues of the covariance operator $\Gamma$ and the eigenvalues of $\mathbf{M}$ is given by
@@
$\phi_{k}(\mathbf{t})=\frac{1}{\sqrt{Nl_{k}}}\sum_{n=1}^{N}[v_{k}]_{n}\left\{X_{n}(\mathbf{t})-\mu(\mathbf{t})\right\},\quad\mathbf{t}\in\mathcal{T},\quad k=1,2,\dots,N,$ (34)
where $[v_{k}]_{n}$ is the $n$th entry of the vector $v_{k}$. The scores are then computed as the inner-product between the multivariate curves and the multivariate eigenfunctions and are given by
$\mathfrak{c}_{nk}=\sqrt{Nl_{k}}[v_{k}]_{n},\quad n=1,2,\dots,N,\quad k=1,2,\dots,N.$ (35)
These results can be extended in a natural way if all the curves are expanded in a general basis of functions, defined in Equation (7). The derivation of these equalities, as well as the derivation of the eigencomponents using the expansion of the curves in a general basis of function and the Gram matrix, are given in Appendix B in a slighty more general framework where the observation weights are not equal.
@@ 4.3 Computational complexity
We describe the time complexity for the computation of the MFPCA algorithm using the covariance operator and the inner product matrix. Considering the observation of $N$ curves with $P$ features, we assume that all observations of the feature $p$ are sampled on a common grid of $M_{p}$ points. For $a\in\mathbb{N}$, let $M^{a}=\sum_{p}M_{p}^{a}$. Let $K$ be the number of multivariate eigenfunctions to estimate. For the estimation of the eigencomponents using the covariance operator, we have $K\leq K_{+}$. While $K$ has the same interpretation for both the eigendecomposition of the covariance operator and the eigendecomposition of the inner product matrix, in the latter case, it is not computed as the summation over the univariate elements, but rather as the number of components needed to achieve a certain amount of variance explained. Here, we also assume that the curves are perfectly observed, and thus no smoothing step is included in the expression of the time complexity. Note that the smoothing step will often have the same impact on complexity between the approaches.
To estimate the time complexity of an algorithm, we count the number of elementary operations performed, considering a fixed execution time for each. Worst-case time complexity is considered. We first give the time complexity for the estimation of the eigencomponents using the covariance operator by explaining the time complexity of each individual step (see Happ and Greven, (2018) and Section 4.1). For each feature $p$, the time complexity of the estimation of the covariance matrix is $\operatorname{\mathcal{O}}(NM_{p}^{2})$, of the eigendecomposition of the matrix is $\operatorname{\mathcal{O}}(M_{p}^{3})$ and of the univariate score is $\operatorname{\mathcal{O}}(NM_{p}K_{p})$. Therefore, the total time complexity is the sum over the $p$ univariate time complexities. The covariance matrix $\mathbf{Z}$ of the stacked univariate scores $\mathcal{Z}$ is then computed with a time complexity of $\operatorname{\mathcal{O}}(NK_{+}^{2})$, because the dimension of the matrix $\mathcal{Z}$ is $N\times K_{+}$. The eigendecomposition of the matrix $\mathbf{Z}$ has a time complexity of $\operatorname{\mathcal{O}}(K_{+}^{3})$. The final step is to compute the multivariate eigenfunctions and scores. For the estimation of the $K\leq K_{+}$ multivariate eigenfunctions, the time complexity is $\operatorname{\mathcal{O}}(K\sum_{p}M_{p}K_{p})$ and for the estimation of the scores, the time complexity is $\operatorname{\mathcal{O}}(NK^{2})$. Gathering all the results, the final complexity of the estimation of the eigencomponents using the eigendecomposition of the covariance operator is
@@
The number of components $K$ to estimate is usually small compared to the number of curves $N$ or to the total number of sampling points $M^{1}$. Both time complexities can then be reduced to $\mathcal{O}(NM^{2}+M^{3})$ for the diagonalization of the covariance operator and to $\mathcal{O}(N^{2}M^{1}+N^{3})$ using the Gram matrix. If the number of observations is large compared to the total number of sampling points, it thus seems preferable to use the covariance operator to estimate the eigencomponents, while if the total number of sampling points is large compare to the number of observations, the use of the Gram matrix seems better. Note that the number of features $P$ does not have much impact on the computational complexity. These results are confirmed in the simulation (see Section 5.2).
@@ 5 Empirical analysis
Using simulated data, we compare the estimation of the eigencomponents using the diagonalization of the covariance operator and the Gram matrix. The diagonalization of the covariance operator is performed using the methodology of Happ and Greven, (2018). As this methodology is based on the expansion of each univariate feature into univariate principal components, we used univariate FPCA, if the curves are unidimensional, and the Functional Canonical Polyadic-Tensor Power Algorithm (FCP-TPA) for regularized tensor decomposition (Allen,, 2013), if the curves are two-dimensional. We choose the FCP-TPA as it is used by Happ and Greven, (2018) in their algorithm and implemented in their software (Happ-Kurz,, 2020). Note that we could also use a two-dimensional basis expansion such as penalized tensor splines or discrete cosine transform, but we do not investigate these expansions here, as we do not want to prespecify a basis of functions.
The results of the simulation are compared using computation times (CT), the integrated squared error (ISE) risk function for the multivariate eigenfunctions, the log-absolute error ($\log-$AE) risk function for the eigenvalues and the mean integreated squared error (MISE) risk function for the reconstructed data. Let $\phi_{k}$ be the true eigenfunction and $\widehat{\phi}_{k}$ the estimated eigenfunction defined on $\mathcal{T}$. We then define the ISE as
@@
Let $\lambda=\{\lambda_{1},\dots,\lambda_{K}\}$ be the set of true eigenvalues and $\widehat{\lambda}=\{\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{K}\}$ be the set of estimated eigenvalues. We then define the $\log-$AE as
$\log-\text{AE}(\lambda_{k},\widehat{\lambda}_{k})=\log(\lvert\lambda_{k}-\widehat{\lambda}_{k}\rvert),\quad k=1,\dots,K.$ (39)
@@
$\text{MISE}(\mathcal{X},\widehat{\mathcal{X}})=\frac{1}{N}\sum_{n=1}^{N}\left|\!\left|\!\left|X_{n}-\widehat{X}_{n}\right|\!\right|\!\right|^{2}=\frac{1}{N}\sum_{n=1}^{N}\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left\{X_{n}^{(p)}(t_{p})-\widehat{X}_{n}^{(p)}(t_{p})\right\}^{2}{\rm d}t_{p}.$ (40)
Each integral is approximated by the trapezoidal rule with an equidistant grid. We let $\widehat{\phi}$, $\widehat{\lambda}$ and $\widehat{\mathcal{X}}$ be the estimators obtained using the Gram matrix and $\widetilde{\phi}$, $\widetilde{\lambda}$ and $\widetilde{\mathcal{X}}$ the estimators obtained using the covariance operator. For each simulation, we compute the ratios
$\frac{\text{ISE}(\phi_{k},\widehat{\phi}_{k})}{\text{ISE}(\phi_{k},\widetilde{\phi}_{k})},\quad\frac{\log-\text{AE}(\lambda_{k},\widehat{\lambda}_{k})}{\log-\text{AE}(\lambda_{k},\widetilde{\lambda}_{k})},\quad k=1,\dots,K,\quad\text{and}\quad\frac{\text{MISE}(\mathcal{X},\widehat{\mathcal{X}})}{\text{MISE}(\mathcal{X},\widetilde{\mathcal{X}})},$ (41)
and compare them to $1$.

### v2

@@ 3.4 On centering and reducing
For conducting an MFPCA, the features are usually assumed centred (Happ and Greven,, 2018). Prothero et al., (2023) give a complete overview of centering in the context of FDA. Here, we comment on the geometric interpretation of centering in this context and compare with the multivariate scalar case. We focus on the usual centering in FDA, namely $X_{n}^{(p)}(t_{p})-\mu^{(p)}(t_{p}),~{}t_{p}\in\mathcal{T}_{p}$ (referred to object centering in Prothero et al., (2023)). The geometric interpretation of object centering is the same if we refer to the observation space $\mathcal{H}$ or the feature space $\mathbb{R}^{N}$. Within the space $\mathcal{H}$ (resp. $\mathbb{R}^{N}$), centering is interpreted as translating the centre of gravity of the clouds, $\mathrm{G}_{\!\mu}$ (resp. $\mathsf{G}_{\!\mu}$), to the origin point $\mathrm{O}_{\!\mathcal{H}}$ (resp. $\mathsf{O}_{\!\mathbb{R}}$) of the space $\mathcal{H}$ (resp. $\mathbb{R}^{N}$). This transformation, being a translation, does not change the shape of the cloud $\mathcal{C}_{\!P}$ (resp. $\mathcal{C}_{\!N}$). The interpretation is the same as for the centering in the multivariate scalar data context within their observation space.
Concerning the standardization of the data, there are two main proposals in the literature. Happ and Greven, (2018) propose to consider the weights
@@
This standardization is coherent with the derivation of the total inertia of the observation space using the usual distance in $\mathcal{H}$. Chiou et al., (2014) and Jacques and Preda, (2014) propose the weighting function
$w_{p}(t_{p})=\left(\operatorname{Var}X^{(p)}(t_{p})\right)^{-1},\quad t_{p}\in\mathcal{T}_{p}.$ (30)
This corresponds to a standardization of the curves by the standard deviation of the component at each sampling point. The standard deviation curve is estimated as the square root of the diagonal of the covariance function estimates, obtained using a local linear smoother of the pooled data. For each functional feature $p$, this standardization mimics the standardization used for principal components analysis if the number of (scalar) features is infinite. Considering the duality diagram and the total inertia of the clouds with respect to the distance $d_{\Gamma}$, we propose to use the weights
$w_{p}=\left(\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}\right)^{-1}.$ (31)
The total inertia of $\mathcal{C}_{\!P}$ (and $\mathcal{C}_{\!N}$) will thus be equal to the number of components. In every case, we may consider the rescaled elements $\widetilde{X}^{(p)}=w_{p}^{1/2}\{X^{(p)}-\mu^{(p)}\}$ in place of the elements $X^{(p)}$.
@@ 4 Multivariate functional principal components analysis
where $\mathfrak{c}_{k}=\langle\!\langle X-\mu,\phi_{k}\rangle\!\rangle$ are the projections of the centered curves onto the eigenfunctions. We have that $\mathbb{E}(\mathfrak{c}_{k})=0$, $\mathbb{E}(\mathfrak{c}_{k}^{2})=\lambda_{k}$ and $\mathbb{E}(\mathfrak{c}_{k}\mathfrak{c}_{k^{\prime}})=0$ for $k\neq k^{\prime}$. Note that the coefficients $\mathfrak{c}_{k}$ are scalar random variables while the multivariate functions $\phi_{k}$ are vectors of functions. Let us call $\Phi$ the multivariate functional principal component analysis basis. In practice, we use a truncated version of the Karhunen-Loève expansion (33) as the eigenvalues $\lambda_{k}$, and hence the contribution of $\mathfrak{c}_{k}$ to (33), becomes negligible as $k$ goes to infinity. Let
@@
be the truncated Karhunen-Loève expansion of the $p$th feature of the process $X$. For each $p$, the function $\mu^{(p)}$ is the $p$th feature of the multivariate mean function $\mu$ and the set $\{\varphi^{(p)}_{k}\}_{1\leq k\leq K_{p}}$ is a basis of univariate functions in $\mathcal{L}^{2}\left(\mathcal{T}_{p}\right)$, whose elements are not the components of the multivariate functions $\phi_{k}$. In (35), the coefficients $\mathfrak{c}^{(p)}_{k}$ are the projection of the centered curve $X^{(p)}$ onto the eigenfunctions $\varphi_{k}^{(p)}$ and are not (directly) related to the coefficients $\mathfrak{c}_{k}$ in (34).
@@ 4.1 Diagonalization of the covariance operator
We recall here how to estimate the eigencomponents. Following Happ and Greven, (2018, Prop. 5), the multivariate components for $X$ are estimated by a weighted combination of the univariate components computed from each $X^{(p)}$. First, we perform a univariate FPCA on each of the features of $X$ separately. For a feature $X^{(p)}$, the eigenfunctions and eigenvectors are computed using a matrix decomposition of the estimated covariance $C_{p,p}$ from (13). This results in a set of eigenfunctions $\{\varphi_{k}^{(p)}\}_{1\leq k\leq K_{p}}$ associated with a set of eigenvalues $\{\lambda_{k}^{(p)}\}_{1\leq k\leq K_{p}}$ for a given truncation integer $K_{p}$. Then, the univariate scores for a realization $X_{n}^{(p)}$ of $X^{(p)}$ are given by $\mathbf{c}_{nk}^{(p)}=\left\langle X_{n}^{(p)},\varphi_{k}^{(p)}\right\rangle,~{}1\leq k\leq K_{p}$. These scores might be estimated by numerical integration for example. Considering $K_{+}\coloneqq\sum_{p=1}^{P}K_{p}$, we then define the matrix $\mathcal{Z}\in\mathbb{R}^{N\times K_{+}}$, where on each row we concatenate the scores obtained for the $P$ features of the $n$th observation: $(\mathbf{c}_{n1}^{(1)},\ldots,\mathbf{c}_{nK_{1}}^{(1)},\ldots,\mathbf{c}_{n1}^{(P)},\ldots,\mathbf{c}_{nK_{P}}^{(P)})$. An estimation of the covariance of the matrix $\mathcal{Z}$ is given by $\mathbf{Z}=(N-1)^{-1}\mathcal{Z}^{\top}\mathcal{Z}$. An eigenanalysis of the matrix $\mathbf{Z}$ is carried out to estimate the eigenvectors $\boldsymbol{v}_{k}$ and eigenvalues $\lambda_{k}$. Finally, the multivariate eigenfunctions are estimated as a linear combination of the univariate eigenfunctions using
@@
$\mathfrak{c}_{nk}=\mathcal{Z}_{{n,\cdot}}\boldsymbol{v}_{k},\quad 1\leq n\leq N,\quad 1\leq k\leq K_{+},$
where $\mathcal{Z}_{{n,\cdot}}$ is the $n$th row of the matrix $\mathcal{Z}$. We refer the reader to Happ and Greven, (2018) for the derivation of the eigencomponents of the covariance operator if the curves are expanded in a general basis of functions.
@@ 4.2 Diagonalization of the inner product matrix
We use the duality relation between row and column spaces of a data matrix to estimate the eigencomponents of the covariance operator. Consider the inner-product matrix $\mathbf{M}$, with entries defined in (5) and assuming that all observations are equally weighted, i.e., for all $n=1,\dots,N$, $\pi_{n}=1/N$. Let $\{l_{k}\}_{1\leq k\leq N}$ such that $l_{1}\geq\dots\geq l_{N}\geq 0$ be the set of eigenvalues and $\{\boldsymbol{u}_{k}\}_{1\leq k\leq N}$ be the set of eigenvectors of the matrix $\mathbf{M}$. The relationship between all nonzero eigenvalues of the covariance operator $\Gamma$ and the eigenvalues of $\mathbf{M}$ is given by
@@
$\phi_{k}(\mathbf{t})=\frac{1}{\sqrt{Nl_{k}}}\sum_{n=1}^{N}[\boldsymbol{u}_{k}]_{n}\left\{X_{n}(\mathbf{t})-\mu(\mathbf{t})\right\},\quad\mathbf{t}\in\mathcal{T},\quad k=1,2,\dots,N,$ (37)
where $[\boldsymbol{u}_{k}]_{n}$ is the $n$th entry of the vector $\boldsymbol{u}_{k}$. The scores are then computed as the inner-product between the multivariate curves and the multivariate eigenfunctions and are given by
$\mathfrak{c}_{nk}=\sqrt{Nl_{k}}[\boldsymbol{u}_{k}]_{n},\quad n=1,2,\dots,N,\quad k=1,2,\dots,N.$ (38)
The derivations of these equalities are given in Appendix B in a slighty more general framework where the observation weights $\pi_{n}$ are not equal. These results can be extended in a natural way if all the curves are expanded in a general basis of functions, see Section LABEL:sub:with_a_basis_expansion in the Supplementary Material.
@@ 4.3 Computational complexity
We describe the time complexity for the computation of the MFPCA algorithm using the covariance operator and the Gram matrix. Considering the observation of $N$ curves with $P$ features, we assume that all observations of feature $p$ are sampled on a common grid of $M_{p}$ points. For $a\in\mathbb{N}$, let $M^{a}=\sum_{p}M_{p}^{a}$. Let $K$ be the number of multivariate eigenfunctions to estimate. For the estimation of the eigencomponents using the covariance operator, we have $K\leq K_{+}$. While $K$ has the same interpretation for both the eigendecomposition of the covariance operator and the eigendecomposition of the inner product matrix, in the latter case, it is not computed as the summation over the univariate elements, but rather as the number of components needed to achieve a certain amount of variance explained. Here, we also assume that the curves are perfectly observed, and thus no smoothing step is included in the expression of the time complexity. Note that the smoothing step will often have the same impact on complexity between the approaches as the smoothing is a preprocessing step.
To estimate the time complexity, we count the number of elementary operations performed, considering a fixed execution time for each. Worst-case time complexity is considered. We first give the time complexity for the estimation of the eigencomponents using the covariance operator by explaining the time complexity of each individual step (see Happ and Greven, (2018) and Section 4.1). For each feature $p$, the time complexity of the estimation of the covariance matrix is $\operatorname{\mathcal{O}}(NM_{p}^{2})$, of the eigendecomposition of the matrix is $\operatorname{\mathcal{O}}(M_{p}^{3})$ and of the univariate score is $\operatorname{\mathcal{O}}(NM_{p}K_{p})$. Therefore, the total time complexity is the sum over the $p$ univariate time complexities. The covariance matrix $\mathbf{Z}$ of the stacked univariate scores $\mathcal{Z}$ is then computed with a time complexity of $\operatorname{\mathcal{O}}(NK_{+}^{2})$, because the dimension of the matrix $\mathcal{Z}$ is $N\times K_{+}$. The eigendecomposition of the matrix $\mathbf{Z}$ has a time complexity of $\operatorname{\mathcal{O}}(K_{+}^{3})$. The final step is to compute the multivariate eigenfunctions and scores. For the estimation of the $K\leq K_{+}$ multivariate eigenfunctions, the time complexity is $\operatorname{\mathcal{O}}(K\sum_{p}M_{p}K_{p})$ and for the estimation of the scores, the time complexity is $\operatorname{\mathcal{O}}(NK^{2})$. Gathering all the results, the final complexity of the estimation of the eigencomponents using the eigendecomposition of the covariance operator is
@@
The number of components $K$ to estimate is usually small compared to the number of curves $N$ or to the total number of sampling points $M^{1}$. Both time complexities can then be reduced to $\mathcal{O}(NM^{2}+M^{3})$ for the diagonalization of the covariance operator and to $\mathcal{O}(N^{2}M^{1}+N^{3})$ using the Gram matrix. If the number of observations is large compared to the total number of sampling points, it thus seems preferable to use the covariance operator to estimate the eigencomponents, while if the total number of sampling points is large compare to the number of observations, the use of the Gram matrix seems better. Note that the number of features $P$ does not have much impact on the computational complexity, in the sense that the important part is the total number of sampling points. One component with $1000$ sampling points will have the same computational complexity as $100$ components with $10$ sampling points. These results are confirmed in the simulation (see Section 5.2).
@@ 5 Empirical analysis
Using simulated data, we compare the estimation of the eigencomponents and the reconstruction of the curves using the diagonalization of the covariance operator, (a) based on univariate FPCA as well as (b) P-splines basis expansion, and (c) the diagonalization of the Gram matrix. These methods will be refered as (a) (Tensor) PCA, (b) 2D/1D B-Splines and (c) Gram respectively. For the diagonalization of the covariance operator, we consider the methodology of Happ and Greven, (2018). In the case of MFPCA based on the expansion of each univariate feature into univariate principal components ((Tensor) PCA), the eigencomposition of the 1-dimensional curves is performed using univariate FPCA and the eigendecomposition of the 2-dimensional curves is calculated using the Functional Canonical Polyadic-Tensor Power Algorithm (FCP-TPA) for regularized tensor decomposition (Allen,, 2013). In the case of MFPCA based on P-splines basis expansion (2D/1D B-Splines), data are expanded in B-splines basis with suitable smoothness penalty (Eilers and Marx,, 1996).
The results of the simulation are compared using computation times (CT), the integrated squared error (ISE) for the multivariate eigenfunctions, the relative squared error (RSE) for the eigenvalues and the mean relative squared error (MRSE) for the reconstructed data. Let $\phi_{k}$ be the true eigenfunction and $\widehat{\phi}_{k}$ the estimated eigenfunction defined on $\mathcal{T}$. We then define the ISE as
@@
Let $\lambda=\{\lambda_{1},\dots,\lambda_{K}\}$ be the set of true eigenvalues and $\widehat{\lambda}=\{\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{K}\}$ be the set of estimated eigenvalues. We then define the RSE as
$\text{RSE}(\lambda_{k},\widehat{\lambda}_{k})=\left(\lambda_{k}-\widehat{\lambda}_{k}\right)^{2}/\lambda_{k}^{2},\quad k=1,\dots,K.$ (42)
@@
$\text{MRSE}(\mathcal{X},\widehat{\mathcal{X}})=\frac{1}{N}\sum_{n=1}^{N}\left|\!\left|\!\left|X_{n}-\widehat{X}_{n}\right|\!\right|\!\right|^{2}=\frac{1}{N}\sum_{n=1}^{N}\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left\{X_{n}^{(p)}(t_{p})-\widehat{X}_{n}^{(p)}(t_{p})\right\}^{2}{\rm d}t_{p}.$ (43)
Each integral is approximated by the trapezoidal rule with an equidistant grid.